RYAN HAS THE FOLLOWING DATA:

VC/UNIT

$54.40

VC PERCENTAGE 80%

FIXED COSTS $114,240

HOW MANY UNITS MUST RYAN SELL TO BREAK EVEN?

UNITS___________

2.

REFER TO QUESTION 1. HOW MANY UNITS MUST RYAN SELL TO ACHIEVE AFTER

TAX NET INCOME (ATNI) = $79,968? (ASSUME 30% TAX RATE)

UNITS__________

3.

JANE HAS A W/A CM = $8.25. ATNI IS CURRENTLY $109,725. WHAT WILL ATNI BE IF

JANE IS ABLE TO SELL 1,400 UNITS ABOVE CURRENT LEVELS? (ASSUME 30% TAX

RATE)

ATNI__________.

4.

TOM HAS THE FOLLOWING DATA FOR PRODUCTS A, B AND C:

PRODUCT A CM/UNIT = $9, MIX PERCENTAGE = 25%

PRODUCT B CM/UNIT = $5, MIX PERCENTAGE = 60%

PRODUCT C CM/UNIT = $3, MIX PERCENTAGE = 15%

FIXED COSTS = $79,800. WHAT IS THE TOTAL $CM OF PRODUCT A WHEN TOM’S

ATNI = $35,055. (ASSUME TAX RATE = 25%)

PRODUCT A $CM___________
Answer #1

1.
Contribution Margin Ratio = 100% - 80% = 20%
Contribution Margin per unit = $54.40/80x20 = $13.60

Break Even Units = Fixed Costs / Contribution margin per unit
= $114240/13.60 = 8400 units

2.
Income before tax = $79968/(1-0.30) = $114240
Units required = (Fixed Costs+Target Income) / Contribution margin per unit
= ($114240+114240)/13.60 = 16800 units

3.
ATNI = $109725 + 1400 x 8.25 x 0.70 = $117810

4.
CM = Fixed Costs + ATNI/(1-t)
= $79800 + 35055/(1-.025) = $126540

Product A CM = $126540 x 25% = $31635